what is the formula mass of calcium sulfate dihydrate


1
Expert's answer
2021-01-29T08:26:44-0500

Calcium Sulfate dihydrate = "CaSO_4.2H_2O"


formula mass = Ca + S + 4×O + 2(2×H + O) = 40 + 32 + (16×4) + 2(1×2 + 16) = 172g.